OpenStudy (anonymous):

A sample of gas occupies 313mL at STP. What volume in L will the gas occupy at 88C and 501torr?

OpenStudy (joannablackwelder):

Use PV=nRT to calculate the number of moles of gas you have. Then use that number of moles with the new temp and pressure to find the new volume. Be careful to use the units consistent with your choice of R.

OpenStudy (zpupster):

here is the formula i use for this P1V1/T1=P2V2/T2 STP= 273.15° Kelvin and 760 torr. so plugging in our values 88C=361.15K and 313 ml=.313L (760 torr)(.313L)/(273.15° K)=(501 torr)(V2 in ?L)/(361.15K)